# Eliminating Backdoors in Neural Code Models via Trigger Inversion

### v1

@@ 6. Conclustion
In this paper, we propose EliBadCode, a novel backdoor defense technique based on trigger inversion. EliBadCode aimed at eliminating backdoors in NCMs, ensuring secure code understanding. By PL-specific trigger vocabulary generation and sample-specific trigger position identification, EliBadCode reduces the search space for trigger optimization and minimizes the impact of adversarial perturbations, respectively. Our experiments show that EliBadCode can effectively invert the trigger of the given backdoored NCM. Through trigger unlearning, EliBadCode can reduce the average ASR of backdoored NCMs to a minimum of 0.24% without impacting their performance on clean inputs.

### v2

@@ 7. Conclusion
In this paper, we propose EliBadCode, a novel backdoor elimination technique for ensuring secure code understanding. By PL-specific trigger vocabulary generation and sample-specific trigger position identification, EliBadCode reduces the search space for trigger optimization and minimizes the impact of non-backdoor perturbations, respectively. Our experiments show that EliBadCode can effectively invert the trigger of the given backdoored NCM. Through trigger unlearning, EliBadCode can reduce the average ASR of backdoored NCMs to a minimum of 0.24% without impacting their performance on normal inputs.
8. Data Availability
Our source code and experimental data are available at (Sun et al., 2025).
Acknowledgement
We would like to thank anonymous reviewers for their insightful comments. This work is supported partially by the National Research Foundation, Singapore, and DSO National Laboratories under the AI Singapore Programme (AISG Award No: AISG2-GC-2023-008), the National Natural Science Foundation of China (61932012, 62372228, U24A20337), the Fundamental Research Funds for the Central Universities (14380029), the Open Project of State Key Laboratory for Novel Software Technology at Nanjing University (Grant No. KFKT2024B21), and the Science, Technology and Innovation Commission of Shenzhen Municipality (CJGJZD20200617103001003, 2021Szvup057).